Find the expected count and the contribution to the chi-square statistic for the (Group 1, No) cell in the two way table below.

Yes No

Group 1 710 266

Group 2 1166 323

Round your answer for the expected count to one decimal place, and your answer for the contribution to the chi-square statistic to three decimal places.

Expected count=

contribution to the chi-square statistic=
Answer #1

here expected count for the (Group 1, No) cell =row total*coluimn total/grand total =976*589/2465=233.2

contribution to the chi-square statistic for the (Group 1, No) cell =(266-233.2)2/233.2 =4.613   (please try 4.610 if this comes wrong)